Alex is 15 years older than Cristina. In 4 years, the sum of their ages will be 91. How old are they now?
Answer by ikleyn(52787)

.
A = C + 15            (1)    ("Alex is 15 years older than Cristina.")
(A+4) + (C+4) = 91    (2)    ("In 4 years, the sum of their ages will be 91")


Substitute expression A = C+15  from eq(1) into eq(2). You will get


((C+15)+4) + (C+4) = 91


2C + 15 + 4 + 4 = 91

2C = 91 - 15 - 4 - 4 = 68  ====>

C = 68/2 = 34.


Answer.  C is 34 years old.  A is 34+15 = 49 years old.  


Check.   (34+4) + (49+4) = 38 + 53 = 91.   ! Correct !

Solved.

-------------------

There is a bunch of lessons on age word problems
    - Age problems and their solutions
    - A fresh formulation of a traditional age problem
    - Really intricate age word problem
    - Selected age word problems from the archive
    - Age problems for mental solution
in this site.

Read them and become an expert in solving age problems.

Also, you have this free of charge online textbook in ALGEBRA-I in this site
    - ALGEBRA-I - YOUR ONLINE TEXTBOOK.

The referred lessons are the part of this online textbook under the topic "Age word problems".


Save the link to this online textbook together with its description

Free of charge online textbook in ALGEBRA-I
https://www.algebra.com/algebra/homework/quadratic/lessons/ALGEBRA-I-YOUR-ONLINE-TEXTBOOK.lesson

to your archive and use it when it is needed.